Gabby07:

A rectangular field is 130 yards long and 85 yards wide. Give the length and width of another rectangular field that has the same perimeter but a larger area.

Florisalreadytaken:

oopsies, lapsus again.... -- you are asked to find a rectangle with the same perimeter, but a LARGER area. to get a rectangle with the same perimeter, but a larger area, we have to take something off the length, and and add the same reduced ammount to the width of it, so you will end up with a vertically-longer rectangle, and dimmed on the sides.

1 attachment
Florisalreadytaken:

the width has got to be < than 130. but > than 85. lets take an example, which you can use as an answer. \( 115 \) long and \( 100 \) wide \(P_1=170+260 \Rightarrow P_1=430\) \(P_2=200+230 \Rightarrow P_2=430\) \(A_1=130 \times 85 \Rightarrow A_1=11050\) \(A_2=115 \times 100 \Rightarrow A_2=11500\) you can see that there is a difference, and out rectangle is BIGGER.

surjithayer:

@surjithayer wrote:
Perimeter =2(130+85)=2 \times 215=430 yards. Let the dimensions of new field be x and y. then 2(x+y)=430 x+y=430/2=215 y=215-x \[area~ A=xy=x(215-x)=215x-x^2\] \[\frac{ dA}{dx }=215-2x\] \[\frac{ dA}{ dx }=0,\] gives 215-2x=0 2x=215 x=215/2=107.5 \[\frac{ d^2A }{ dx^2}=-2 \] at x=107.5,\[\frac{ d^2A}{ dx^2 }=-2<0\] Hence A is maximum at x=107.5 yards and y=210-107.5=107.5yards.

surjithayer:

the rectangle has max. area if it is a square of each side=107.5 yards